A chemist performed this reaction and obtained 45.5 g BaSO4. How much HCI

was also produced?

Molar masses: 233.4 g/mol BaSO4,

36.5 g/mol HCI

A. 9.49 g HCI

C. 5.35 g HCI

B. 388 g HCI

D. 7.12 g HCI


1
Expert's answer
2022-06-03T08:06:02-0400

Answer D: 7.12 g HCl


m (HCl) = m (BaSO4) / M (BaSO4) × M (HCl) = 45.5/233.4×36.5= 7.12 g